hi
i need your help on this one age problem that im stumped on
grandma is three years younger than grandpa and at their 30th anniversery, grandma was 17/18 of grandpa's age. how old were they when they got married?
thanks for your help

Answer by stanbon(75887)

grandma is three years younger than grandpa and at their 30th anniversery, grandma was 17/18 of grandpa's age. how old were they when they got married?
--------------------------------------
Let grandpa's age be "x" ; grandma's age is "x-3"
------------------------------
30th anniversary ages:
x-3 = (17/18)x
(1/18)x = 3
x = 54 (grandpa's age on 30th anniversary)
x-30 = 24 grandpa's age when he married.
Grandma was 3 years younger or 21 when she married.
-----------------------------------
Cheers,
Stan H.